# SocialEyes: Scaling mobile eye-tracking to multi-person social settings

### v3

@@ 1. Introduction
Tracking eye movements has mostly been limited to single-person settings under controlled environments due to the immobility of traditional eye-tracking hardware and the complexities of unrestricted multi-person data collection and analysis. With the recent development of affordable, mobile eye-tracking technology, it is increasingly possible to conduct eye-tracking “in the wild” (see e.g. (Fasold et al., 2021; Kulke and Hinrichs, 2021; Reitstätter et al., 2020; Saxena et al., 2023; Vandemoortele et al., 2018; Fridman et al., 2018)). For single-person use cases, mobile eye-tracking glasses are most commonly applied to collect data in unrestricted settings but a number of challenges remain unsolved for their adoption in multi-person settings. For instance, these glasses record individual gaze from the egocentric perspective of the wearer; data from all differing perspectives of wearers need to be aligned when conducting multi-person studies. To effectively study collective gaze dynamics in naturalistic social settings, it is crucial to develop automated ways to synchronise and analyse the eye-tracking data of multiple individuals.
@@
A schematic of our approach to solve the problem of multi-person eye-tracking in such collective viewing experiences is detailed in Figure 1. We imagine a scenario where multiple people observe a shared scene from different viewing angles. The eye movements of each viewer are recorded independently from their egocentric perspective (egoview) using commercial-grade eye-tracking glasses. However, to study collective gaze patterns, the wearer-specific gaze data needs to be synchronised and mapped onto a shared coordinate space (semantic gaze mapping). We record a centralview of the shared scene from a camera placed behind all observers. We employ Network Time Protocol (NTP) to temporally synchronise data from all devices. The synchronised gaze data from all viewers is then mapped to the shared coordinate space i.e. the centralview by finding robust feature matches between egoview–centralview image pairs and performing a projective transformation. Following our approach we design a software framework–SocialEyes. We implement and test SocialEyes during two public events in which we recorded eye data from 30 participants simultaneously. We detail our choices of eye-tracking hardware, the underlying infrastructure, and the implementation specifics of the framework. In validating our approach, we provide visualisation and analysis metrics for multi-person eye-tracking data. Overall, we make the following contributions in this paper:
@@
Introduce and implement standard and new visualisation and analysis techniques for multi-person eye-tracking data. We provide solutions for visualising both egocentric and homography-transformed gaze data and heatmaps, with methods to visualise all individually or in parallel. The transformed gaze heatmaps offer particular promise in analysing group attention and perception. Similarly, analysis metrics like convex hull area may offer novel insights into collective gaze behaviour.
Problem setting
@@ 2. Related Work
More scalable and automated attempts at unrestricted eye-tracking are actively being made under the umbrella field of ‘pervasive eye-tracking’ (Kasneci, 2017) and motivated by advances in portable head-mounted eye trackers that have recently become more affordable and accurate. The increased availability of open-sourced and automated analysis tools (Panetta et al., 2019; Li et al., 2006) has boosted the adoption and potential applications of these mobile eye trackers, which are increasingly being employed in unrestricted natural environments with static scenes, such as art galleries (Reitstätter et al., 2020). However, to date, a synchronised multi-person eye-tracking study in dynamic real world social settings has not been done and consequently the analysis methods to evaluate such large-scale data are not readily available. To solve these challenges we design and implement SocialEyes.
@@ 3. SocialEyes: Design Concept
System Design
@@ 3.1.2. Streaming
The streaming mode is designed for real-time applications where data are collected and processed concurrently. Processing data in real-time is desirable for interactive applications such as gaze-contingent interfaces. This mode of operation demands sophisticated hardware and software optimisations to achieve good performance. Streaming of data in this mode is facilitated by the open-source event streaming platform–Apache Kafka–providing high-throughput and low-latency. Data streams are processed by Kafka topics, each comprising of multiple partitions that store data in background Java programs called brokers while ensuring load balancing and scalability. Finally, separate consumer modules read these data streams for a) calculating homography matrix, b) writing to local storage, and/or c) publishing the data to a dedicated Flask server for real-time monitoring.
@@ 3.2.2. GlassesRecord
The GlassesRecord module is engaged in the recording mode to send recording start and stop signals to each pair of eye-tracking glasses. The module features an intuitive and reliable operator interface to efficiently trigger and monitor actions on multiple devices simultaneously (see example implementation in 4). Actions can be triggered selectively on a subset of devices to ensure fault tolerance, for example, the ability to restart recording on one pair of glasses if it fails, rather than having to restart on all devices. The module also provides functionality to annotate recordings with custom hand-marked events.
@@ 3.2.6. Homography
The homography module implements semantic gaze mapping (as defined in our approach 1) from egoview coordinate space to centralview coordinate space simultaneously for multiple devices. Applying a pin-hole camera model, the egoview captured from each wearer’s perspective can be assumed to lie on a planar surface at an infinite distance. The projective transformation between the egoview and centralview for the same scene is then mathematically related by a planar homography. Incoming gaze samples (x, y coordinates) and egoview frames from each device are paired together with the nearest centralview frame using the marked timestamps from the respective input modules of each data stream. The module accommodates varying data sampling rates of the three data streams, since each stream’s sampling resolution depends upon factors like hardware configuration, recording parameters, and network bandwidth. The final temporal resolution of the homography module output is the same as the sampling rate of the provided centralview (i.e., for each centralview frame, the homography module attempts to map all gaze points from the provided egoviews+gaze pairs to that centralview frame). This mapping is accomplished by finding robust feature point correspondences between each egoview-centralview image pair. Matched key points are then used to calculate a homography matrix that is finally used to transform gaze from egoview to centralview coordinate space.
The performance of this module depends significantly on detecting and tracking robust feature point correspondences in dynamic, real-world social settings which often involve challenging artefacts caused by illumination changes, motion distortion, low-textured regions, and occlusions. In recording mode, the module can utilize large deep learning models, such as the state-of-the-art SuperGlue (Sarlin et al., 2020) model demonstrated in our Utility Test (see 4), thanks to the absence of computational and time constraints. However, for real-time operation in streaming mode, simpler algorithms like Bay et al. (2008); Rublee et al. (2011); Tomasi and Kanade (1991) are preferred, balancing the trade-off between speed and accuracy.
@@ 3.2.7. Visualisation
Single-person gaze and egoview: Overlays gaze synchronised to the egoview stream for each selected viewer as a separate video file with an optional overlay of associated timestamps for each data sample.
Multi-person gaze and egoview: Renders a single video with all selected viewers’ egoviews with the corresponding gaze points overlay (as in the previous option) in a 2D grid of customisable size.
Single-person egocentric and transformed gaze view: Stacks the time synchronised egoview and centralview frames for a single viewer along with the corresponding gaze and transformed gaze overlay on each view respectively.
Multi-person transformed gaze on centralview: Presents the centralview with all time synchronised gaze coordinates from the selected viewers overlaid on top. The transformed gaze for each viewer can be plotted either as customisable circles containing a unique ID and outline colour, or as a heatmap overlay calculated from the 2D distribution of all transformed gaze points.
Multi-person egocentric views with transformed gaze on centralview: Combines the egoviews from all selected viewers and plots it as a grid with the time synchronised centralview in the centre cell. Gaze for each viewer is overlaid in their respective egoview and the centre cell displays all the transformed gaze points similar to the previous option.
@@ 3.2.8. Analysis
The Analysis module provides methods to preprocess and evaluate the recorded egocentric gaze data as well as the homography-transformed gaze data for collective gaze analysis metrics. We design the followign metrics including traditional eye-tracking measures, such as heatmap similarity and correlation, scaled up to multi-person use cases, and novel collective measures, such as convex hull area, to quantify gaze dynamics in social group settings:
Average Gaze Velocity: Gaze velocity is calculated as a time series of inter-sample Euclidean distances of gaze (x, y coordinates) samples. The average gaze velocity over a given timeframe quantifies mean precision of the gaze predictions which reflects overall reliability of the gaze predictions.
Heatmap: Gaze heatmaps are calculated from a series of x and y gaze coordinates by convolving a 2D Gaussian filter with sigma (standard deviation of the gaussian kernel) equal to the approximate size of the fovea (i.e., one degree of visual angle).
Heatmap Similarity (SIM) and Correlation (CC): SIM is calculated as the histogram intersection between the distributions of two heatmaps (ranges between 0 and 1 where 1 represents identical distributions) and CC is represented by the Pearson correlation coefficient between them to quantify the degree of linear relationship between their pixel intensities.
Stationary Gaze Entropy: Stationary gaze entropy is calculated by organising gaze coordinates in a given time frame into spatial bins of roughly one degree of visual angle (after conversion to the same coordinate space). Shannon’s entropy equation is then applied to this discrete probability distribution of gaze to calculate the average level of uncertainty or the overall predictability of the gaze coordinates. The obtained entropy quantifies gaze dispersion in the given time frame with a higher entropy or uncertainty indicating a wider distribution of gaze across the visual field (Holmqvist et al., 2011).
Convex Hull Area: Convex Hull Area provides a dynamic measure of gaze dispersion over time. It is calculated as the area of the smallest convex polygon that encloses all provided gaze samples at a given time point. The convex hull area was initially adapted to eye-tracking analysis by Goldberg& Kotval (1999).
@@ 4.1.1. Apparatus
Eye-tracking equipment: We used the Pupil Labs Neon eye-tracking devices (Baumann and Dierkes, 2023) that apply a deep learning pipeline to provide reliable gaze predictions in unrestricted settings. At the core of these eye trackers is a hardware module equipped with binocular infrared cameras that capture eye images with a resolution of 192 x 192 pixels at 200 Hz and a front-facing RGB camera that records the egoview with a resolution of 1600 x 1200 pixels at 30 Hz. The fixed geometry of eye and egoview cameras on the module eliminates the requirement of explicit eye-egoview calibration before each recording; the gaze is recorded in egoview coordinates. The module is mounted to a 3D-printed, wearable frame that connects to an accompanying Android smartphone (Motorola Edge 40 Pro). The minimal form factor of the frames is designed to be unobtrusive and effortless for participants, making them a good fit for longer-duration studies. The accompanying Android phone runs a proprietary vendor application–“Neon Companion”–that hosts a suite of HTTP REST Application Programming Interfaces (APIs) enabling remote streaming of data and controlling operations such as start/stop recording. During the data collection event, each pair of eye-tracking glasses was connected to the companion smartphone device running Pupil’s Labs NEON Companion android application in the background. The phones were locked and connected to an Anker 6-in-1 USB-C hub that extended the ports to connect them to ethernet, power, and the NEON glasses concurrently. Each smartphone was secured in a bag attached to the respective participant’s chair, ensuring free movement of the seated participant (see Figure 3B). The ethernet cable connected the device to a network switch placed under the false floor, in the centre of the audience rows. The global NTP server for all smartphones, the central camera, and the server computers were all set to the University’s NTP server accessible on the local network.
@@
Centralview camera: A PTZOptics pt30x-sdi-g2 camera was mounted central to the stage, positioned behind the audience seating. The camera captured an unobstructed view of the stage with a resolution of 720x480 px at 60 Hz. The camera feed was streamed on the network via RTSP with H264 encoding. Figure 3A demonstrates an example frame from the central camera.
Presentation screen: The film was presented on a 14.5 ft (384.5 cm x 216 cm) Samsung LH015IER LED cabinet with a resolution of 1920x1080 pixels at a 60 Hz refresh rate and a pixel pitch of 1.5 mm.
Computing servers: The GlassesRecord and CentralCam modules were executed on separate desktop servers running Ubuntu 18.20 LTS Linux distribution. Note that both modules have low CPU requirements and could be easily executed on a single general-purpose server, as long as sufficient storage is available for long-duration recordings. We leveraged the modular structure of our framework to run the modules on separate servers to allow separate operators and monitors for the two modules.
Network Infrastructure: Each pair of Neon glasses requires a bandwidth of 6-8 Megabytes per second (MBps) to stream gaze and egoview data. For real-time streaming operations, it is critical to consider the available bandwidth on a network to identify how many concurrent eye-tracking glasses’ data can be streamed. In this test, we used the recording mode of operation which has lower bandwidth requirements, since real-time data streams do not need to be accessed on the server. Nonetheless, we leveraged the stationary positioning of participants in our setup to provide precise timing information and optimal network reliability by connecting each eye-tracking device to a network switch (TP Link TL-SG1048) via ethernet. The network did not have internet access, to disable any background network activity on the connected Android smartphones.
Storage: egoview videos and gaze data were recorded and stored locally on each smartphone through the Neon companion app. The recorded data was later uploaded to Pupil Cloud for post-processing of the recorded raw data and then downloaded for further analysis. Time offsets and logs from the GlassesRecord module were stored locally on the server running the module. Centralview video and metrics recorded by the CentralCam module were stored locally on the server running the CentralCam module.
@@ 4.1.2. SocialEyes implementation in Recording mode
GlassesRecord: We implemented the GlassesRecord module utilising the pupil-labs-real-time-api python library (https://pupil-labs-realtime-api.readthedocs.io) to send start, stop, discard, and save recording-triggers to each Neon device. Periodic clock offset measurements were recorded every ten seconds. Additionally, since the accompanying smartphones use Android as operating system, we used Android Debug Bridge (ADB) to retrieve status metrics including battery level, storage level, network ping, ADB server status, and connected USB devices. The status metrics facilitated continuous monitoring of devices and recording status. ADB was also used to remotely control the Android smartphones for remote troubleshooting and verifying indicators of failed or corrupted recordings that were not reported by the vendor application. These operations were important to prevent data loss and ensure that we would never need to interrupt the live social event to access a participants’ phone in the event of an error. The operator interface for this module was implemented as a Terminal User Interface (TUI) built using the python Textual library (https://textual.textualize.io/), providing low runtime requirements and cross platform operation. The TUI displays a list of all eye-tracking devices on the local network, along with colour-coded status metrics for each device (battery/storage level, network ping, ADB server status, recording status, etc.). Multiple devices can be selected from this list to perform a desirable action with dedicated keystrokes. The following actions can be performed on all selected devices using the TUI: Start Recording; Save Recording (end and save recordings locally); Cancel Recording (end and discard ongoing recordings); Restart App (use ADB to restart the Neon Companion app, critical for troubleshooting app crashes remotely); Reconnect ADB (reconnect to wireless ADB, critical for recovering lost connection to devices).
CentralCam: The CentralCam module was implemented with OpenCV (Bradski, 2008) and FFMPEG (https://ffm peg.org) libraries to retrieve the RTSP stream from the central camera. The incoming stream was compressed to a MPEG4 container using an H264 codec. A tolerance of 600 dropped frames was added to ignore intermittent packet drops on the network before the module terminates recording. FPS was calculated over a window of 180 frames and stored with other metrics to a CSV file format.
Homography, Analysis, and Visualisation: We applied a pre-trained Graph Neural Network–-SuperPoint/ SuperGlue (Sarlin et al., 2020)–to find robust feature point correspondences in our Homography module. Blinks were identified using the Neon blink-detection algorithm (Labs, 2023) and filtered out of the raw gaze data before homography, analysis and visualisation. The operator interface for these modules was implemented as a command-line utility using the Python questionary (questionary.readthedocs.io) library. The command-line utility offered formatting of the directory structure for compatibility with modules functions, selecting a subset of recordings to perform functions, such as homography, visualisation, gaze metrics calc‘ulation, etc., and storing output to the local filesystem. Gaze heatmaps were generated using the viridis scientific colormap (van der Walt et al., 2020), which is designed for accurate perceptual uniformity and accessibility for viewers with color vision deficiencies.
Visual angle calculation: Visual angle conversion was done by estimating the distances of each of the three participant rows from the videowall. The distances were 758 cm for the first, 885 cm for the second, and 1012 cm for the third participant row. Since the physical dimensions of the video wall were 384.5 cm for width and 216 cm for height, the diagonal (441.02 cm) projected a visual angle of 32.4° for the first, 28° for the second, and 25.6° for the third row. From the egoview recordings of participants in each row, the diagonal of the videowall was estimated to be 515 pixels for the first, 441 pixels for the second, and 396 pixels for the third row. The visual angle to pixel ratio was therefore calculated to be approximately 16 (i.e. one visual degree was represented by roughly 16 pixels in the 1600x1200 resolution egoview recordings).

### v4

@@ 1. Introduction
Tracking eye movements has mostly been limited to single-person settings under controlled environments due to the immobility of traditional eye-tracking hardware and the complexities of unrestricted multi-person data collection and analysis. With the recent development of affordable, mobile eye-tracking technology, it is increasingly possible to conduct eye-tracking “in the wild” (see e.g., (Fasold et al., 2021; Kulke and Hinrichs, 2021; Reitstätter et al., 2020; Saxena et al., 2024; Vandemoortele et al., 2018; Fridman et al., 2018)). For single-person use cases, mobile eye-tracking glasses are most commonly applied to collect data in unrestricted settings but a number of challenges remain unsolved for their adoption in multi-person settings. For instance, these glasses record individual gaze from the egocentric perspective of the wearer; data from all differing perspectives of wearers need to be aligned when conducting multi-person studies. To effectively study collective gaze dynamics in naturalistic social settings, it is crucial to develop automated ways to synchronise and analyse the eye-tracking data of multiple individuals.
@@
A schematic of our approach to solve the problem of multi-person eye-tracking in such collective viewing experiences is detailed in Figure 1. We imagine a scenario where multiple people observe a shared scene from different viewing angles. The eye movements of each viewer are recorded independently from their egocentric perspective (egoview) using commercial-grade eye-tracking glasses. However, to study collective gaze patterns, the wearer-specific gaze data need to be synchronised and mapped onto a shared coordinate space (semantic gaze mapping). We record a centralview of the shared scene from a camera placed behind all observers. We employ Network Time Protocol (NTP) to temporally synchronise data from all devices. The synchronised gaze data from all viewers is then mapped to the shared coordinate space i.e. the centralview by finding robust feature matches between egoview–centralview image pairs and performing a projective transformation. Following our approach we design a software framework–SocialEyes. We implement and test SocialEyes during two public events in which we recorded eye data from 30 participants simultaneously. We detail our choices of eye-tracking hardware, the underlying infrastructure, and the implementation specifics of the framework. In validating our approach, we provide visualisation and analysis metrics for multi-person eye-tracking data. Overall, we make the following contributions in this paper:
@@
Introduce and implement standard and new visualisation and analysis techniques for multi-person eye-tracking data. We provide solutions for visualising both egocentric and homography-transformed gaze data and heatmaps, with methods to visualise all individually or in parallel. The transformed gaze heatmaps offer particular promise in analysing group attention and perception. Similarly, analysis metrics like normalised contour area may offer novel insights into collective gaze behaviour.
[Problem setting]Figure 1. Fully described in the text.
@@ 2. Related Work
More scalable and automated tools for unrestricted eye-tracking are actively being developed under the umbrella field of ‘pervasive eye-tracking’ (Kasneci, 2017). These are motivated by advances in portable head-mounted eye trackers that have recently become more affordable and accurate. The increased availability of open-sourced and automated analysis tools (Panetta et al., 2019; Li et al., 2006) has boosted the adoption and potential applications of these mobile eye trackers, which are increasingly being employed in unrestricted natural environments with static scenes, such as art galleries (Reitstätter et al., 2020). However, to date, a synchronised multi-person eye-tracking study in dynamic real world social settings has not been done and consequently the analysis methods to evaluate such large-scale data are not readily available. To solve these challenges we design and implement SocialEyes.
@@ 3. SocialEyes: Design Concept
[System Design]Figure 2. In recording mode, gaze and egoview data are captured by the GlassesRecord module, while the CentralCam module records centralview data. The data is stored locally or in the cloud for retrieval by the Analysis, Visualization, and Homography modules. In streaming mode, the GlassesStream module records glasses data, which is transmitted via Apache Kafka to subsequent modules.
@@ 3.1.2. Streaming
The streaming mode is designed for real-time applications where data are collected and processed concurrently. Processing data in real-time is desirable for interactive applications such as gaze-contingent interfaces. This mode of operation demands sophisticated hardware and software optimisations to achieve good performance. Streaming of data in this mode is facilitated by the open-source event streaming platform–Apache Kafka–providing high-throughput and low-latency. Data streams are processed by Kafka topics, each comprising of multiple partitions that store data in background Java programs called brokers while ensuring load balancing and scalability. Finally, separate consumer modules read these data streams for a) calculating homography matrices, b) writing to local storage, and/or c) publishing the data to a dedicated Flask server for real-time monitoring.
@@ 3.2.2. GlassesRecord
The GlassesRecord module is engaged in the recording mode to send recording start and stop signals to each pair of eye-tracking glasses. The module features an intuitive and reliable operator interface to efficiently trigger and monitor actions on multiple devices simultaneously (see example implementation in Section 4). Actions can be triggered selectively on a subset of devices to ensure fault tolerance, for example, the ability to restart recording on one pair of glasses if it fails, rather than having to restart on all devices. The module also provides functionality to annotate recordings with custom hand-marked events.
@@ 3.2.6. Homography
The homography module implements semantic gaze mapping from egoview coordinate space to centralview coordinate space simultaneously for multiple devices (see Figure 1). Applying a pin-hole camera model, the egoview captured from each wearer’s perspective can be assumed to lie on a planar surface at an infinite distance. The projective transformation between the egoview and centralview for the same scene is then mathematically related by a planar homography. Incoming gaze samples (x, y coordinates) and egoview frames from each device are paired together with the nearest centralview frame using the marked timestamps from the respective input modules of each data stream. The module accommodates varying data sampling rates of the three data streams, since each stream’s sampling resolution depends upon factors like hardware configuration, recording parameters, and network bandwidth. The final temporal resolution of the homography module output is the same as the sampling rate of the provided centralview (i.e., for each centralview frame, the homography module attempts to map all gaze points from the provided egoviews+gaze pairs to that centralview frame). This mapping is accomplished by finding robust feature point correspondences between each egoview-centralview image pair. Matched key points are then used to calculate a homography matrix that is finally used to transform gaze from egoview to centralview coordinate space.
The performance of this module depends significantly on detecting and tracking robust feature point correspondences in dynamic, real-world social settings which often involve challenging artefacts caused by illumination changes, motion distortion, low-textured regions, and occlusions. In recording mode, the module can utilize large deep learning models, such as the state-of-the-art SuperGlue (Sarlin et al., 2020) model demonstrated in our Utility Test (see Section 4), thanks to the absence of computational and time constraints. However, for real-time operation in streaming mode, simpler algorithms like Bay et al. (2008); Rublee et al. (2011); Tomasi and Kanade (1991) are preferred, balancing the trade-off between speed and accuracy.
@@ 3.2.7. Visualisation
[…]
@@ 3.2.8. Analysis
The Analysis module provides methods to preprocess and evaluate the recorded egocentric gaze data as well as the homography-transformed gaze data for collective gaze analysis. It includes the following metrics combining traditional eye-tracking measures, such as heatmap similarity and correlation, scaled up to multi-person use cases, with novel collective measures, such as normalised contour area, to quantify gaze dynamics in social group settings:
[…]
@@ 4.1.1. Apparatus
Eye-tracking equipment: We used the Pupil Labs Neon eye - tracking devices (Baumann and Dierkes, 2023) that apply a deep learning pipeline to provide reliable gaze predictions in unrestricted settings. At the core of these eye trackers is a hardware module equipped with binocular infrared cameras that capture eye images with a resolution of 192 x 192 pixels at 200 Hz and a front-facing RGB camera that records the egoview with a resolution of 1600 x 1200 pixels at 30 Hz. The fixed geometry of eye and egoview cameras on the module eliminates the requirement of explicit eye-egoview calibration before each recording; the gaze is recorded in egoview coordinates. The module is mounted to a 3D-printed, wearable frame that connects to an accompanying Android smartphone (Motorola Edge 40 Pro). The minimal form factor of the frames is designed to be unobtrusive and effortless for participants, making them a good fit for longer-duration studies. The accompanying Android phone runs a proprietary vendor application–“Neon Companion” (v2.8.2) –that hosts a suite of HTTP REST Application Programming Interfaces (APIs) enabling remote streaming of data and controlling operations such as start/stop recording. During the data collection event, each pair of eye-tracking glasses was connected to the companion smartphone device running Pupil’s Labs NEON Companion android application in the background. The phones were locked and connected to an Anker 6-in-1 USB-C hub that extended the ports to connect them to ethernet, power, and the NEON glasses concurrently. Each smartphone was secured in a bag attached to the respective participant’s chair, ensuring free movement of the seated participant (see Figure 3B). The ethernet cable connected the device to a network switch placed under the false floor, in the centre of the audience rows. The global NTP server for all smartphones, the central camera, and the server computers were all set to the University’s NTP server accessible on the local network.
[Study setup]Figure 3. Fully described in the text
@@ 4.1.2. SocialEyes implementation in Recording mode
GlassesRecord: We implemented the GlassesRecord module utilising the pupil-labs-real-time-api python library (https://pupil-labs-realtime-api.readthedocs.io) to send start, stop, discard, and save recording triggers to each Neon device. Periodic clock offset measurements were recorded every ten seconds. Additionally, since the accompanying smartphones use Android as operating system, we used Android Debug Bridge (ADB) to retrieve status metrics including battery level, storage level, network ping, ADB server status, and connected USB devices. The status metrics facilitated continuous monitoring of devices and recording status. ADB was also used to remotely control the Android smartphones for remote troubleshooting and verifying indicators of failed or corrupted recordings that were not reported by the vendor application. These operations were important to prevent data loss and ensure that we would never need to interrupt the live social event to access a participants’ phone in the event of an error. The operator interface for this module was implemented as a Terminal User Interface (TUI) built using the python Textual library (https://textual.textualize.io/), providing low runtime requirements and cross platform operation. The TUI displays a list of all eye-tracking devices on the local network, along with colour-coded status metrics for each device (battery/storage level, network ping, ADB server status, recording status, etc.). Multiple devices can be selected from this list to perform a desirable action with dedicated keystrokes. The following actions can be performed on all selected devices using the TUI: Start Recording; Save Recording (end and save recordings locally); Cancel Recording (end and discard ongoing recordings); Restart App (use ADB to restart the Neon Companion app, critical for troubleshooting app crashes remotely); Reconnect ADB (reconnect to wireless ADB, critical for recovering lost connection to devices).
[…]